OpenStudy (anonymous):

What is the slope of the line between (3, –4) and (–2, 1)?

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2-y1)/(x2-x1) m = (1-( -4))/(-2-3) m = (1+4)/(-2-3) m = 5/(-5) m = -1

OpenStudy (anonymous):

Thank you, that helped a lot! (:

jimthompson5910 (jim_thompson5910):

you're welcome